OpenStudy (anonymous):

how do you solve? x-2/3=6 1/3

OpenStudy (anonymous):

this one is simple you will simply: x-2/3 = 6 1/3 add 2/3 to both sides to get x =6 3/3 which is 7 so x=7

OpenStudy (anonymous):

thank you !

OpenStudy (anonymous):

you are welcome i hope this helped explain it to you so you can do them in the future on your own instead of me just giving you the answer